# Joint Distributed Generation Maximization and Radial Distribution Network Reconfiguration

### v1

@@ III-D Comparing Mixed Integer Conic Programming Relaxation
1 1.1025 25 7.7518 0.39754 $-4.9991$ 0.092606
2 1.0964 0.2645 0 0 0.50264 $-0.19736$
Conic relaxation in (13)
[…]
1 1.1025 25 7.9991 0.64489 $-4.9991$ 0.092606
2 1.0915 25 0 0 0.75 0.049999
With the problem data in Fig. 4 and the variable values in Table V, we can verify that the solutions of (12) and (13) satisfy all constraints in the respective problems. For example, the reactive power balance equation in (1b) for bus 1 for (12) is $q^{G}_{1}-q^{L}_{1}=0.39754-0.5=-0.10246=(-1)\times 0.092606+(-0.19736)+0.0075\times 25=-Q_{1}+Q_{2}+x_{01}\ell_{1}$. The analogous equation holds for the solution of (13) as well: $q^{G}_{1}-q^{L}_{1}=0.64489-0.5=0.14489=(-1)\times 0.092606+0.049999+0.0075\times 25=-Q_{1}+Q_{2}+x_{01}\ell_{1}$. However, while equation (5d) at bus 2 holds for the solution of (12): $\nu_{1}\ell_{2}=1.1025\times 0.2645=0.2916=0.50264^{2}+{(-0.19736)}^{2}=P_{2}^{2}+Q_{2}^{2}$, (13b) remains an inequality instead of equality for the solution of (13): $\nu_{1}\ell_{2}=1.1025\times 25=27.5625>0.565=0.75^{2}+0.049999^{2}=P_{2}^{2}+Q_{2}^{2}$. Indeed, a posterior load flow analysis based on $(p^{G},q^{G})$ from (13) reveals that the bus voltages are 1, 1.0539 and 1.0511 per unit, respectively. Similarly, the currents for the lines are 522.53 A and 51.235 A. The voltage and current upper limits are both violated by the solution of conic relaxation. The aforementioned comparison is repeated for all benchmarks in Table I. The conic relaxation solutions fail to satisfy equality (5d) in all cases, and result in voltage and current violations.
The objective functions in (12) and (13) have a significant impact on the tightness of conic relaxation. For instance, if we instead minimize the total line loss (as in [2, 16]), then the solutions of (13) also solve (12) for all benchmark instances in Table I, except for the instances not achieving convergence within the time limit of one hour. Following [28], we consider the trade-off between DG maximization and line loss minimization by minimizing in (12) and (13) the following weighted sum objective:
where $0\leq\rho\leq 1$ is an algorithmic parameter for trade-off. With $\rho$ increasing from 0 to 1, the minimization of (14) transitions from DG maximization to loss minimization. By solving (12) and (13) with (14) and running a posterior load flow analysis based on the resulting DG set points and network configuration, the trade-off can be visualized. Fig. 5 shows the Pareto curves for the 33bw benchmark with universal line rating of 600 A and $K=6$.
Figure 5: Pareto curves of 33bw benchmark (600A line rating) due to the proposed model in (12) and conic relaxation in (13)
For $\rho>0.5$, the two Pareto curves coincide since the conic relaxation solution satisfies (5d). However, when $\rho\leq 0.5$ conic relaxation fails, and conic relaxation no longer depicts the true trade-off. The extra “DG output” of conic relaxation is paid by violating the voltage and current limits. We note that the part of the Pareto curve of the proposed method for $\rho\leq 0.5$ is in fact not obtained by solving (12) with (14) but rather by maximizing DG as in (12) with an additional constraint upper bounding total line loss.
III-E Comparison with Linearized DistFlow Approximation
We evaluate a variant of (12) by replacing the DistFlow equations in (5) with its linear approximation (i.e., LinDistFlow equations [1]). The LinDistFlow variant ignores the loss-related terms multiplying $\ell_{n}$ in (5a), (5b) and (5c). Also, equation (5d) is dropped. Substituting the DistFlow equations with the LinDistFlow equations changes (12) to a mixed integer linear program, which is easier to solve. However, the voltage variables in the LinDistFlow variant overestimate the true voltages (e.g., [29, eq. (2a) and (3)]). This may lead to false voltage lower limit satisfaction. Further, since the squared current term $\ell$ is missing it is not possible to directly impose current upper limit constraints. Indeed, for our benchmark instances severe violation of current upper limit is experienced. To counter this, a surrogate current upper limit constraint can be imposed, similar to (5d), as
In (15) the product in the left-hand-side can be simplified using the linearization technique illustrated in (6). Below we call LinDistFlow the variant of (12) with LinDistFlow equations and (15). The inclusion of (15) relieves some undervoltage and overcurrent issues. However, this is not enough. For instance, for the 33bw benchmark with 600 A line rating and $K=2$, the actual voltages by a posterior load flow analysis and the voltages perceived to be true by the LinDistFlow optimization model (i.e., $(\nu)^{1/2}$) are very different as indicated by Fig. 6.
Also, current upper limits are violated in multiple lines (worst case violation being 115% of rating) despite the surrogate constraint in (15). For LinDistFlow with $K=4$ the result is even more extreme. The load flow analysis by Matpower fails to converge. In contrast, solving (12) for the 33bw 600 A benchmark with $K=2$ and $K=4$, respectively, leads to constraint-abiding control with guaranteed optimality.
III-F Comparison with Equivalent Formulation by Jabr et. al.
We compare the proposed method with the reconfiguration formulation adopted from [2] by changing the objective from total line loss minimization to DG maximization. Almost all constraints in [2] are linear except
$2u^{(n,k)}u^{(k,n)}\geq R_{nk}^{2}+T_{nk}^{2},\quad\text{$\{n,k\}$ is a line}$ (16)
In (16), all symbols are continuous decision variables and hence (16) are rotated second order cone inequalities. In fact, (16) is a relaxation. The model in [2] is equivalent to (12) only when (16) hold as equalities. Similar to conic relaxation discussed in Section III-D, when minimizing a “standard” objective such as total line loss the inequalities in (16) are often satisfied as equalities at optimality. However, for DG maximization (16) remain inequalities and the resulting control violates voltage and current limits. For example, for the 33bw benchmark with 600 A line rating, solving the formulation in [2] with $K=4$ leads to voltage violation up to 1.25 pu and current violation up to 137% of line rating. To rectify the situation, we can impose (16) as equalities. However, imposing (16) as equalities renders the formulation in [2] very difficult to solve. For instance, for 33bw with a 300-second time limit Gurobi terminates with optimality gap of 256%, 287%, 342% and 205% for $K=2,4,6,8$, respectively (cf. Table II).
IV Case Study - Real Network
IV-A 533-bus Distribution System Description
In this section, the ROPF problem (12) is applied to a real medium voltage distribution system in southern Sweden: the 533-bus network described in [30]. With the expansion of distributed energy resources (DERs), many DSOs wish to assess to what extent DERs can be integrated in their systems without violating operational limits. This is referred to as performing a hosting capacity (HC) analysis. In [31], four operational limits that constrain the HC in distribution networks are outlined – voltage, overloading, protection and harmonics. With our proposed method we are imposing voltage and current limits, thereby dealing with the first two. Protection is partly included by requiring radial operation, but no further protection limits (i.e. short-circuit current) are set up. The system is assumed to be three-phase balanced without overtones, thus harmonics mitigation is not considered. Within these limits, the HC can be assessed and enhanced, using e.g. network reconfiguration, and our proposed ROPF method does this jointly.
The network in [30] operates at 12 kV with a smaller part at 135 kV, covers about 20$\times$30 km and serves about 30,000 inhabitants plus an industrial area. It is connected to the regional 135 kV grid at a single feed-in station, modeled as a slack bus. A system overview is seen in Fig. 7.
@@ IV Case Study - Real Network
The existing wind and solar DG capacity amounts to 34.5 MW, installed at 262 nodes with capacities ranging from 3 kW to 2 MW per node. To avoid curtailment of already existing DG in the system, production from existing DG is embedded in $p^{L}_{n}$ by subtracting hourly generation from hourly load to obtain the net load data for each node. In the year 2022, the total net load in the system varied from 44.6 MW in Dec 16, 08:00 to -4.8 MW in Jul 10, 14:00. The latter will be referred to as the minimum net load hour.
The voltage constraints in (8) are set to $\sqrt{\underline{\nu_{n}}}=0.95$ pu and $\sqrt{\overline{\nu_{n}}}=1.05$ pu, whereas the current constraints $\overline{\ell_{kn}}$ in (9) are set individually for each line, according to their respective rated currents. All system data was provided by the DSO. In the real network, there are often several line segments and conductor types connecting two substations, while in the model these are aggregated to single lines with equivalent rated currents and impedances. A full description of this aggregation procedure can be found in [30]. Once the aggregation has been made, the single lines have rated phase currents ranging from 93-1200 A, phase resistances from 0.003-1.86 $\Omega$ and phase reactances from 0.001-0.73 $\Omega$.
Furthermore, the real network is a three-phase system, whereas the optimization algorithm is designed for single-phase. Since the system is assumed to be three-phase balanced, this was managed by dividing all loads and generation capacities by a factor three, effectively treating each phase as a separate system. After simulations were run, the actual power flows were obtained by multiplying all power values post-optimization with a factor three.
The final Matpower-model of the 533-bus system can be found at github.com/MATPOWER/matpower/blob/master/data/case533mt_lo.m with net load values from the minimum net load hour.
IV-B Reconfiguration with 533-bus Example
The ROPF problem in (12) is solved for the 533-bus system through two separate case studies. The first study involves the construction of a multiple node generation scenario, representing a potential future expansion of wind power. The second study is a spatially comprehensive DG expansion site analysis, where additional DG capacity is added only to a single node in the system at the time, but iterating through the entire system, to see how high the HC is in each specific location and how much it can be increased through network reconfiguration.
@@ IV-B Reconfiguration with 533-bus Example
IV-B1 Wind scenario
@@
Solving (12) for this generation setup, $\sum_{n}p^{G}_{n}=17.6$, $19.2$ and $19.3$ MW for $K=0$, $2$ and $4$ respectively. As an initial observation, the added DG capacity of $21\times 2=42$ MW is significantly curtailed both with and without reconfiguration. With $K=2$, the HC increases by 1.6 MW (or 9.1%) compared to the normal topology. This increase is due to a pair of switch events that divides a cluster of four units in the southeast on two separate feeder lines, in contrast to the normal topology where they share a common feeder that turns out to be a bottleneck (see Fig. 9 and Fig. 10). If four switch events are allowed ($K=4$), the algorithm returns the same first pair of switch events as with $K=2$, but also another pair of switch events closer to the distribution station LF. This has a more limited effect, increasing HC by only 0.1 MW from $K=2$.
Figure 9: Wind scenario in the normal radial topology ($K=0$), zoomed in at the southeastern part of the network.
Figure 10: Wind scenario after reconfiguration ($K=2$),
the black arrows indicate where switch events occurred.
IV-B2 DG expansion site analysis
In the DG expansion site analysis, additional DG capacity is added to a single node one at a time. This added capacity is large enough for the apparent power rating of the unit to not be an active constraint in (7b), allowing us to identify the HC of the grid. After adding DG to a single node, (12) is solved with $K=0$, $2$ and $4$, to enable a comparison of the HC before and after reconfiguration. This procedure is repeated for all 532 non-slack nodes in the system, sometimes referred to as the iterative or detailed method [32]. No single optimization took more than 10 mins, which is considered reasonable for a site analysis.
In Fig. 11, the results from the $532\times 3$ optimizations in the DG expansion site analysis can be seen. The $x,y$-value indicates the location of a node in the network and the bar height is the HC at that node in MW. The blue part of each bar is the HC in the normal radial topology, whereas the red and yellow part is the additional HC including reconfiguration with $K=2$ and $4$ respectively.
Figure 11: 3D bar chart displaying the HC at each node. The blue, red and yellow parts are with $K=0$, $2$ and $4$ respectively.
Tall bars in Fig. 11 indicate a high HC, typically found at distribution substations or large industry nodes. On average, the HC is also higher at urban nodes than at rural. A large share of red or yellow in a bar indicates that the HC at that node can be improved substantially by reconfiguration. A large share of the urban nodes, but also some rural, exhibit this substantial increase in HC from reconfiguration.
In Fig. 12, these relative HC improvements from network reconfiguration are presented explicitly. The figure displays how many of the nodes, $x\%$, that exhibit a relative HC increase of $y\%$ or higher from reconfiguration.
@@ IV-B2 DG expansion site analysis
IV-C Discussion of 533-bus Example
The results from this section show that our proposed method can be used to address relevant problems that many DSOs today are facing. The wind scenario shows that the ROPF method can return new, more beneficial topologies for DG expansion. All else equal, our proposed method can thus reduce the need for costly network expansion/reinforcement. The DG expansion site analysis demonstrates how our proposed method can be used to assess and enhance the HC at specific sites. In particular, this can be done with reconfiguration taken into account, adding an additional feature to classic OPF.
Due to local variations and other complicating factors, it is far from certain that the HC is always most limited during the minimum net load hour. The results from this section should thus be interpreted as proof of concept and areas of use of our proposed method, rather than as a comprehensive HC analysis of this particular system.
V Conclusion
DG maximization is more difficult than loss reduction for network reconfiguration as the former objective renders standard techniques such as local optimization, conic relaxation and linearization approximation ineffective, thus raising caution against their over-generalization and highlighting the importance of decision models featuring exact AC power flow equations (or equivalents). With today’s optimization technology, the DG maximization problem can be handled reasonably well for moderately sized networks in a time scale relevant for control center applications. The ability to handle exact AC power flow constraints helps not only in network reconfiguration but also in parameter estimation, fault detection and contingency analysis. However, experiments are needed to evaluate the computation performance in these use cases. On another note, the DistFlow equations appear to be “easier constraints” than other equivalents such as AC power flow equations and the expression in [2]. The effect of problem formulations on branch-and-bound algorithm performance is not well understood. Insights on the effect are highly desirable.

### v2

@@ 5 Practical Case Study - Real Network
The existing wind and solar DG capacity amounts to 34.5 MW, installed at 262 nodes with capacities ranging from 3 kW to 2 MW per node. To avoid curtailment of already existing DG in the system, production from existing DG is embedded in $p^{L}_{n}$ by subtracting hourly generation from hourly load to obtain the net load data for each node. In the year 2022, the total net load in the system varied from 44.6 MW in Dec 16, 08:$0$ to -4.8 MW in Jul 10, 14:$0$. The latter will be referred to as the minimum net load hour.
The voltage constraints in (8) are set to $\sqrt{\underline{\nu_{n}}}=0.95$ pu and $\sqrt{\overline{\nu_{n}}}=1.05$ pu, whereas the current constraints $\overline{\ell_{kn}}$ in (9) are set individually for each line, according to their respective rated currents. All system data was provided by the responsible DSO, Kraftringen, which operates local distribution networks in multiple areas in southern Sweden, including the 533-bus network in question. In the real network, there are often several line segments and conductor types connecting two substations, while in the model these are aggregated to single lines with equivalent rated currents and impedances. A full description of this aggregation procedure, and the model construction overall, can be found in [21]. Once the aggregation has been made, the single lines have rated phase currents ranging from 93-1200 A, phase resistances from 0.003-1.86 $\Omega$ and phase reactances from 0.001-0.73 $\Omega$.
Furthermore, the real network is a three-phase system, whereas the optimization algorithm is designed for single-phase. Since the system is assumed to be three-phase balanced, this was managed by dividing all loads and generation capacities by a factor three, effectively treating each phase as a separate system. After simulations were run, the actual power flows were obtained by multiplying all power values post-optimization with a factor three. The final Matpower-model of the 533-bus system can be found at github.com/MATPOWER/matpower/blob/master/data/case533mt_lo.m with net load values from the minimum net load hour.
5.2 Reconfiguration with 533-bus Example
The ROPF problem in (12) is solved for the 533-bus system through two separate case studies. The first study involves the construction of a multiple node generation scenario, representing a potential future expansion of wind power. The second study is a spatially comprehensive DG expansion site analysis where additional DG capacity is added only to a single node in the system one at a time but iterating through the entire system, to see how high the HC is in each specific location and how much it can be increased through network reconfiguration.
@@ 5.2 Reconfiguration with 533-bus Example
Additionally, in both the wind scenario and the DG expansion site analysis, results are given for $K=0$, 22 and 44 respectively. Naturally, the method can be applied also with higher $K$-values. However, this usually renders quickly diminishing marginal returns as shown in [15], [21] and Sec. 5.2.1. In other words, the highest increases in HC are generally gained from the first switch events. Moreover, the optimization time rapidly increases as the search space grows exponentially with $K$. Finally, DSOs are in practice seldom interested in making too drastic topology changes in their networks based on specific load scenarios. The low $K$-scenarios reveal small topology changes from the normal configuration with potentially high rewards, which is attractive for grid planners.
5.2.1 Wind scenario
@@
Solving (12) for this generation setup, $\sum_{n}p^{G}_{n}=17.6$, $19.2$ and $19.3$ MW for $K=0$, 22 and 44 respectively. As an initial observation, the added DG capacity of $21\times 2=42$ MW is significantly curtailed both with and without reconfiguration. With $K=2$, the HC increases by 1.6 MW (or 9.1%) compared to the normal topology. This increase is due to a pair of switch events that divides a cluster of four units in the southeast on two separate feeder lines, in contrast to the normal topology where they share a common feeder that turns out to be a bottleneck (see Fig. 9). If four switch events are allowed ($K=4$), the algorithm returns the same first pair of switch events as with $K=2$, but also another pair of switch events closer to the distribution station LF. This has a more limited effect, increasing HC by only 0.1 MW from $K=2$.
Figure 9: Left: Wind scenario in the normal radial topology ($K=0$), zoomed in at the southeastern part of the network. Right: Wind scenario after reconfiguration ($K=2$), the black arrows indicate where switch events occurred.
5.2.2 DG expansion site analysis
In the DG expansion site analysis, additional DG capacity is added to a single node one at a time. This added capacity is large enough for the apparent power rating of the unit to not be an active constraint in (7b), allowing us to identify the HC of the grid. After adding DG to a single node, (12) is solved with $K=0$, 22 and 44. This procedure is repeated for all 532 non-slack nodes in the system, sometimes referred to as the iterative or detailed method [31]. No single optimization took more than 10 mins, which is considered reasonable for a site analysis.
In Fig. 10, the results from the $532\times 3$ optimizations in the DG expansion site analysis can be seen. The $x,y$-value indicates the location of a node in the network and the bar height is the HC at that node in MW. The blue part of each bar is the HC in the normal radial topology, whereas the red and yellow part is the additional HC including reconfiguration with $K=2$ and 44 respectively.
Figure 10: 3D bar chart displaying the HC at each node. The blue, red and yellow parts are with $K=0$, 22 and 44 respectively.
Tall bars in Fig. 10 indicate a high HC, typically found at distribution substations or large industry nodes. On average, the HC is also higher at urban nodes than at rural. A large share of red or yellow in a bar indicates that the HC at that node can be improved substantially by reconfiguration. A large share of the urban nodes, but also some rural, exhibit this substantial increase in HC from reconfiguration. In Fig. 11, these relative HC improvements from network reconfiguration are presented explicitly. The figure displays how many of the nodes, $x\%$, that exhibit a relative HC increase of $y\%$ or higher from reconfiguration.
@@ 5.2.2 DG expansion site analysis
6 Conclusion and Future Work
DG maximization is more difficult than loss reduction for ROPF as the former objective renders standard techniques such as local optimization, MICP relaxation and linearization approximation ineffective, thus raising caution against their over-generalization and highlighting the importance of decision models featuring exact AC power flow equations (or equivalents). With a proper choice of problem formulation and optimization software package (e.g., DistFlow equations and Gurobi SBB), DG maximizing ROPF can indeed be handled reasonably well for moderately sized networks in a time scale relevant even for control center applications.
In this paper, the three-phase balanced problem has been considered. This is a natural starting point and a reasonable generalization as many distribution networks across the world are in balanced or close to balanced operation, the networks using three-phase feeders instead of single-phase laterals [22],[25]. For future work however, attention should be paid also to the unbalanced case, which is a reality in many networks and an important source of complexity. The ability of DG inverters and other sources of voltage control to compensate unbalances between phases could then be exploited further. Indeed, there exist variants of the DistFlow equations for three-phase unbalanced and meshed networks. Incorporating these equations in the ROPF model and identifying the proper solution approach are worth further investigation.